OpenStudy (anonymous):

x: 1 over 2 x + 2 = 22.

OpenStudy (anonymous):

(1/2)x + 2 = 22 x + 4 = 44 x = 44 - 4

OpenStudy (anonymous):

1/2x+2=22 1/2x=22-2 1/2x=20 x=10

OpenStudy (anonymous):

@cornitodisc what did you do?

OpenStudy (anonymous):

He divided by 2 instead of multiplying at the end, I was curious too lol.

OpenStudy (anonymous):

yeah,,you right,,i was wrong supposedly x=40